Past experience indicates that the monthly long distance telephone bill per household in a particular community is normally distributed, with a mean of Sh. 1012 and a standard deviation of Sh. 327. After an advertising campaign that encouraged people to make long distance telephone calls more frequently, a random sample of 57 households revealed that the mean monthly long distance bill was Sh. 1098. Can we conclude at the 10% significance level that the advertising campaign was successful


1
Expert's answer
2021-05-25T13:26:07-0400

"\mu = 1012 \\

\sigma= 327 \\

\bar{x}=1098 \\

n=57 \\

H_0: \mu=1012 \\

H_1: \mu > 1012 \\

z=\frac{\bar{x}-\mu}{\sigma/ \sqrt{n}} \\

= \frac{1098-1012}{327/ \sqrt{57}} = 1.985"

The calculated p-value can be calculated as follows.

"P= 1-P(Z<1.985) \\

= 1 -0.9764 = 0.0236"

The calculated p-value is 0.0236. From the p-value, the alternate hypothesis is accepted. We can conclude at the 10% significance level that the advertising campaign was successful.